OpenStudy (abmon98):

The point A has co-ordinates (-1,6) and the point B has coordinates (7,2). Find the equation of the perpendicular bisector of AB. give your answer in the form y=mx+c A point C on the perpendicular bisector has coordinates (p,q). the Distance OC is 2 units where O is the origin. write down the distance OC is 2 units, where O is the origin. write down two equations involving p and q and hence find the coordinates of the possible positions of C

OpenStudy (abmon98):

 x1+x2/2,y1+y2/2)=(3,4) m of perp bisector=2 y=mx+c 4=2(3)+c c=-2 y=2x-2 q=2p-2 O(0,0) C(p,q) ((x2-x1)^2+(y2-y1)^2)^1/2=length between two points ((p-0)^2+(q-0)^2)^1/2=2 (p^2+q^2)=4 p^2+(2p-2)^2=4 p^2+(2p-2)(2p-2)=4 p^2+4p^2-4p-4p+4=4 5p^2-8p=0 p(5p-8) p=0 or 8/5 q=2p-2 plug p=0 q=-2 plug p=8/5 q=2*8/5-2 q=1.2